using D' Alembert method, find the deflection of a vibrating string of unit length having fixed ends, with initial velocity zero and initial deflection f(x)=asin2nx


1
Expert's answer
2022-01-19T08:35:34-0500

Solution:

1-D wave equation can be written as:

"\dfrac{\partial^{2} u}{\partial t^{2}}=c^{2} \dfrac{\partial^{2} u}{\partial x^{2}}"

Here "f(x)=a\sin^2nx"

Using "D^{\prime}" Alembert's method, the solution can be given as:

"y(x,t)=\dfrac12[f(x+ct)+f(x-ct)]
\\=\dfrac12[a\sin^2n(x+ct)+a\sin^2n(x-ct)]
\\=\dfrac a2[\sin^2(nx+nct)+\sin^2(nx-nct)]"

"=\dfrac a2[\dfrac{1-\cos 2(nx+nct)}{2}+\dfrac{1-\cos 2(nx-nct)}{2}]" [Using "\sin^2t=\dfrac{1-\cos 2t}2" ]

"=\dfrac a4[{2-\cos 2(nx+nct)}-\cos 2(nx-nct)]
\\=\dfrac a4[{2-2\cos (\dfrac{2(nx+nct)-2(nx-nct)}{2}}) \cos (\dfrac{2(nx+nct)+2(nx-nct)}{2})]"

"=\dfrac a4[2-2\cos (nct) \cos (nx)]
\\=\dfrac a2[1-\cos (nct) \cos (nx)]"